OpenStudy (anonymous):

Simplify. -5a(2b + 3c) Select one of the options below as your answer: A. -10ab + 3c B. -10ab - 15ac C. 10ab - 15ac D. 10ab + 15ac

OpenStudy (anonymous):

-10ab-15ac

OpenStudy (anonymous):

how did yuhh get dat

OpenStudy (anonymous):

each term in the bracket is multiplied by the term outside: so -5a * 2b = -10ab and -5a * 3c = -15ac so -10ab - 15ac

OpenStudy (anonymous):

i expect u know this already bu ab = a * b and ac = a * c

OpenStudy (anonymous):

ok?